OpenStudy (anonymous):

Explain how you would write the slope-intercept form of an equation when the line passes through points (2, -5) and (0,5).

OpenStudy (anonymous):

m= (y2-y1)/(x2-x1) so, m= (5-(-5))/(0-2) m= (10/-2) m= -5 now use point slope form to write the slope intercept form. y-y1=m(x-x1) y-(-5)=(-5)(x-2) y+5= -5x+10 y=-5x+5 slope intercept form is: y=-5x+5 i'm prettty sure that's right, but you might want to double check it!